What is the probability of obtaining exactly four heads in five flips of a coin, given that at least three are heads? (Enter the probability as a fraction.)
Answer by ewatrrr(24785)

 
Hi,
Independent Events
five flips
P(x = 4Hs|3Hs) Note: last two flips can be HH, HT,TH, TT
P(x = 4Hs|3Hs)= 2/4 = 1/2